OpenStudy (anonymous):

Can anyone help me with a log-log graph? I need to find the formula for a line with points (2, 3.5) and (4550, 75). My text talks about solving the equation simultaneously, but I haven't the foggiest what it means.

OpenStudy (anonymous):

Let (x,y) be a point on the line. Equate the slope of the line using (x,y) and the first given point to the slope calculated using (x,y) and the second given point.\[\frac{y-\frac{35}{10}}{x-2}\text{ = }\frac{y-75}{x-4550}\]Solve the above for y.\[y=\frac{143 }{9096}x+\frac{15775}{4548} \]A plot is attached. You can verify the line equation by plugging the x and y values for the given points into the line equation.

OpenStudy (anonymous):

Thank you, I have to run for now, but I'll plug away at this later and if I get it, I'll hit you off with a medal. Thanks again.

OpenStudy (anonymous):

Yeah, this is not my answer. I need to find a formula in the form y=bx^m

OpenStudy (anonymous):

remember, it's log-log.